A fan delivers 1.75 cu.m/sec of air at a static pressure of 2.5 cm WG through a circular duct with a diameter of 300 mm and discharges it through a 250 - mm diameter duct. If the fan static efficiency is 76%, determine the brake power input to the fan


1
Expert's answer
2021-07-20T21:47:01-0400

the brake power input to the fan :


"BHP = \frac{SP\times CFM}{6356\times \eta_{static}}"


"BHP = \frac{SP\times CFM}{6356\times \eta_{static}}"


Static Pressure (SP) : 2.5  cm WG


CFM – Cubic Feet Per Minute, is the volume of air moved per minute = 3708.03 cfm


Brake Horsepower, is the horsepower absorbed by the fan:


"BHP = \frac{2.5 \times 3708.03}{6356\times 0.76}= 1.92" W